solve by factoring and using the principle of zero products
(x+6)(x-5)=-10

Answer by edjones(8007)

(x+6)(x-5)=-10
x^2+x-30=-10
x^2+x-20=0
(x+5)(x-4)=0
x=-5, x=4
.
Ed